Is it true? The electric field associated with a plane electromagnetic wave is given by E=50 cos(wt-10πx)z bar V/m.its wavelength is 0.5m
1
Expert's answer
2020-03-11T14:20:40-0400

A plane electromagnetic wave with the direction of wave propagation along "x" axis is described by the following equation

(1) "\vec E(t,x)=\hat z E_acos(wt-kx)", where "k=\frac{2\pi}{\lambda}" is a component of wave vector in "x" direction. In our case "k=10\pi", thus "\frac{2\pi}{\lambda}=10\pi" and "\lambda=\frac{1}{5}=0.2 m".

Answer: The task statement is false. It is true "\lambda=0.2m" .